# Is it safe to cross? Interpretable Risk Assessment with GPT-4V for Safety-Aware Street Crossing

### v1

@@ V DISCUSSION
Images Baseline + Optical Flow
There is no car approaching the crosswalk directly, but the pedestrian light is showing a red hand, indicating a stop signal for pedestrians.
There is a vehicle visibly approaching the crosswalk as indicated by the red arrow showing average optical flow towards the crosswalk area.
There is no car approaching the crosswalk directly, but the pedestrian signal is red, indicating a stop sign for pedestrians, which makes it partially dangerous to cross.
There is a vehicle shown in the bottom left image with an arrow indicating its movement toward the crosswalk, which suggests it is unsafe for pedestrians to cross at this moment
TABLE IV: Qualitative results on scene description
@@ V-A The Effect of Visual Knowledge
The analysis reveals that the bounding box feature exhibits relatively low correlation degree and acts as noise when compared to the baseline. This is because not only is the information that can be extracted from object detection already sufficiently extracted from LMM, but the bounding box often obscures important information such as traffic signals, thereby complicating the model’s ability to make accurate judgments.
@@
V-B Does the Chain-of-Thoughts Really Help?
@@ VI CONCLUSIONS
We evaluate a large multimodal model for safety-aware scene understanding in robot navigation, particularly to support blind or low-vision individuals to safely cross streets. By leveraging GPT-4V, we evaluate the capabilities of comprehensive and nuanced understanding of street crossing environments, aiming for development of a trustworthy mobility aid. We share experimental results that provide valuable insights into the effects of various visual features. Optical flow emerged as a significant feature, underscoring the importance of temporal information in enhancing task suitability.
While our findings underscore the effectiveness of the LMM in facilitating safety-aware decisions for street crossing, this is not without its limitations. Notably, the application of the Chain-of-Thought methodology did not yield the anticipated performance improvements. This discrepancy highlights the need for further refinement in prompt engineering and adaptation strategies for visual data. Additionally, our dataset’s sensitivity to specific risk factors and a tendency to overestimate risk in certain scenarios indicate a requirement for greater diversity in data collection.
Future efforts will focus on the impact of different viewpoints on model performance, the importance of temporal understanding in navigation decisions, and the model finetuning based on user-specific preferences. Addressing these areas is essential for advancing our goal of developing a reliable mobility aid to ultimately support BLV peoples’ mobility.

### v2

@@ V-A The Effect of Visual Knowledge
The analysis reveals that the bounding box feature exhibits relatively low correlation degree and acts as noise when compared to the baseline. This is because the information that can be extracted from object detection is already sufficiently captured by the VLM. Moreover, the bounding boxes often obscure important information, such as traffic signals, which complicates the model’s ability to make accurate judgments.
@@
V-B Does Chain-of-Thought Really Help?
@@ VI CONCLUSIONS
We evaluate a vision-language model for safety-aware scene understanding in robot navigation, particularly to support BLV individuals to safely cross streets. By leveraging GPT-4V, we evaluate the capabilities of a comprehensive and nuanced understanding of street crossing environments, aiming for the development of a reliable mobility aid. We share experimental results that provide valuable insights into the effects of various visual features. Optical flow emerged as a significant feature, underscoring the importance of temporal information in enhancing performance.
While our findings underscore the effectiveness of the VLM in facilitating safety-aware decisions for street crossing, this is not without its limitations. Notably, the application of the Chain-of-Thought methodology did not yield the anticipated performance improvements. This discrepancy highlights the need for further refinement in prompt engineering and adaptation strategies for visual data. Furthermore, our dataset’s sensitivity to specific risk factors and its tendency to overestimate risk in certain scenarios highlight the need for greater diversity and quantity in data collection.
Future research will focus on several key areas: investigating the impact of different viewpoints on model performance, exploring the importance of temporal understanding in navigation decisions, comparing the performance of our model with vision-only models, increasing inference speed, and fine-tuning the model based on user-specific preferences. Addressing these areas is crucial for advancing our goal of developing a reliable mobility aid that ultimately supports the mobility of blind and low-vision (BLV) individuals.